OpenStudy (anonymous):

write 10!/(2^5)(5!) as a product of consecutive off integers.

OpenStudy (anonymous):

9X7X5X3X1/ 5X3X1 is it right?

OpenStudy (anonymous):

consecutive odd integers**

OpenStudy (anonymous):

\[is \space \text it \space this? \frac{10!}{2^5\times5!}\]\[or \space \frac{10!5!}{2^5}?\]

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

\[\frac{10 \times 9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1}{2^5 \times 5 \times 4 \times 3 \times 2 \times 1}\]\[\frac{10 \times 9 \times 8 \times 7 \times 6}{2^5}\]\[5 \times 9 \times 7 \times 3\]

OpenStudy (anonymous):

if you don't understand any parts, tell me

OpenStudy (anonymous):

Got it, thank you!